4. Determine the stream function for the following flow field.
Answer #1

=u= x12 1 02 given V=x24 + (x =237 ) ģ ů = uêtres U=x2 , x= x-2xy 2 Y = -(2) 24 w = -(3-2xy) on integrating with respect ay -